# TRABSA: Interpretable Sentiment Analysis of Tweets using Attention-based BiLSTM and Twitter-RoBERTa

### v2

@@ Introduction
This study addresses the need for robust and generalizable SA models by proposing the “Transformer-based Attention Bidirectional LSTM for Sentiment Analysis (TRABSA)" model. The TRABSA model integrates the strengths of transformer-based architectures with attention mechanisms and bidirectional long short-term memory (BiLSTM) networks to enhance the performance and adaptability of SA tasks. Our method seeks to capture the broad diversity of English language usage and offer a more thorough knowledge of sentiment expression in various linguistic situations by combining data from several locations into a single dataset. By using this method, TRABSA can more effectively adjust to the subtle differences in English language usage across various groups, which improves the precision and significance of SA findings. We test the performance of the TRABSA model on a range of DL architectures and datasets, including extensive Twitter and external social media datasets, with a particular emphasis on scalability, accuracy, and consistency. Additionally, we do interpretability assessments utilizing the SHAP and LIME approaches to understand the model’s decision-making mechanism. We show the TRABSA model’s generalizability and robustness through our thorough examination, providing a viable method for SA in various real-world situations.
@@
    We conduct extensive experiments to assess the TRABSA model’s performance on the UK Twitter Dataset, benchmarking against 7 traditional machine learning (ML) models, 4 stacking models, and 4 DL models, demonstrating its superiority and versatility.
@@
The rest of this article is structured as follows: Section “Related Works" reviews state-of-the-art literature in SA using ML-DL and interpretability techniques. In section “Methodology", the data collection and preprocessing techniques, unsupervised text labeling, implemented ML and DL models for benchmarking, architecture and methodology of the proposed TRABSA model are described. The section “Results" presents the experimental setup, evaluation metrics, results, and robustness analysis of the TRABSA model. The SHAP and LIME analysis conducted on the TRABSA model are covered in section “Interpretability Analysis". Section “Disucssions" addresses the findings and implications of our investigation. In conclusion, the article is summarized, and future research directions are outlined in section “Conclusions and Future Directions".
@@
Our proposed methodological framework outlines a structured approach to SA, encompassing several key stages to ensure robustness and effectiveness in model development and evaluation, as shown in Figure 1. The first stage involves gathering relevant data for SA. We extend existing datasets by collecting additional tweets from diverse sources to enhance the dataset’s representativeness and coverage. Following data collection, we perform cleaning and preprocessing tasks to ensure the quality and consistency of the data. This includes removing noise, expanding contractions, handling duplicates, emojis, and missing tweets, and standardizing text formats. The next step involves labeling the data into positive, neutral, and negative sentiments. We leverage the latest updated RoBERTa-based pre-trained transformer model for tokenization and sentiment labeling, enabling accurate and efficient text data processing. The dataset is divided into training, validation, and test sets after it has been labeled. This enables us to use the test set to assess the model’s performance on untested data, refine hyperparameters using the validation set, and train the model on a portion of the data. In this step, we build the SA model based on our suggested hybrid DL architecture. Following model development, we compare the trained model’s performance against baseline models or current state-of-the-art methods. We examine the model’s ability to correctly classify sentiments into positive, neutral, and negative categories using a variety of metrics, including accuracy, precision, recall, and F1-score. Finally, we employ XAI techniques to interpret the model’s predictions and gain insights into its decision-making process. This involves analyzing the model’s internal mechanisms, such as attention weights or feature importance, to understand the factors influencing its predictions and enhance model interpretability.
@@
In this section, we provide a comprehensive overview of the data collection and preprocessing procedures undertaken in our study, which laid the foundation for robust sentiment analysis of tweets.
@@
Extended Datasets: To augment our research’s cross-cultural dimension and overcome the geographic limitations of the benchmark dataset proposed by Qi and Shabrina, we extended the existing UK COVID-19 dataset [16]. The tweets were sourced through a combination of data extraction tools, specifically Twint and the Twitter Academic API. These tools were chosen because of their ability to acquire tweets with geolocation information, which is crucial for conducting geographical analyses. However, it should be noted that only a few 1% of Twitter users actively opt to share their geographic location when composing tweets, and this feature is not enabled by default [40]. The extended datasets, comprising the Global Twitter COVID-19 Dataset and the USA Twitter COVID-19 Dataset, are publicly available in the Extended Covid Twitter Datasets repository [41].
To ensure a comprehensive dataset, we merged the data collected by Twint and the Twitter Academic API. This amalgamation allowed us to access a larger volume of tweets. In identifying tweets related to the COVID-19 pandemic, we employed specific keywords such as “corona" or “covid" in the Twint search configurations and the query field of the Twitter Academic API. This search strategy enabled us to extract tweets and associated hashtags containing these pertinent terms.
@@
    Extended Global COVID-19 Dataset: This extension involved the comprehensive scraping of 411,885 tweets from 32 English-speaking countries. This dataset expansion allowed us to capture sentiment variations across diverse English-speaking regions, as illustrated in Figure 3. In particular, cities such as the “United States," the “United Kingdom," “Australia," and “New Zealand" exhibit high tweet volumes, while several other cities have comparatively lower tweet counts. Figure 4 illustrates word clouds and word frequencies within tweets of the extended datasets. Figure 4 (left) represents a visual summary of the most frequently occurring words in a vast dataset related to the COVID-19 pandemic. At the center of this cloud is the word “covid," which dominates with a staggering 226,463 mentions. Other significant terms around it, such as “vaccine," “case," “test," and “people," indicate the key topics and concerns worldwide during the pandemic. Words like “death," “pandemic," and “health" also hold prominence, highlighting the gravity of public health issues. Additionally, terms like “Trump" and “government" suggest the political dimensions entwined with the pandemic discourse.
@@
    Extended USA COVID-19 Dataset: In addition to the international extension, we further enriched our data by creating an extended dataset focusing exclusively on the United States. This dataset comprised 7,500 tweets meticulously scraped from U.S.-based sources. Including this dataset allows for a closer examination of sentiment dynamics within a specific geographical context. Figure 4 (right) specific to the United States reveals notable trends and sentiments within the country during the COVID-19 pandemic. In this dataset, words like “corona," “coronavirus," and “covid" are prominent, underlining the ubiquitous presence of these terms in American discussions. Interestingly, words like “fool" and “joke" appear, possibly reflecting a spectrum of attitudes towards the pandemic response. Negative expressions like “shit," “fuck," and “die" also emerge, suggesting the emotional intensity and frustration associated with the situation. Terms like “test" and “case" point towards testing and infection rates concerns, while “april" indicates a temporal reference.
Figure 4: This word cloud visualizes the most frequently occurring words in the “Global" (left) and “Only USA" (right) datasets of COVID-19-related tweets. The size of each word corresponds to its frequency in the dataset.
@@
    Removal of Irrelevant Elements: We methodically eliminated any superfluous content that has no bearing on SA, including hashtags (#subject), stated usernames (@username), and any hyperlinks beginning with “www," “http," or “https." We also removed terms that were less than two characters and stop words. Stop words, though common in text, often lack significant sentiment polarity. Despite being often used in texts, stop words frequently lack strong emotive polarity. It’s important to note that negations like “not" and “no" were kept in because removing them may change the sense of whole sentences.
@@
    Handling Repeated Characters: Some users utilize repeating characters in their tweets to highlight intense feelings. Words not present in standard lexicons were transformed into their correct forms to standardize such expressions. For instance, “sooooo goooood" was normalized to “so good."
@@
    Extending Contractions: Removing punctuation after a contraction, such “isn’t" or “don’t," presented difficulties. They were expanded into their full forms to maintain the meaningfulness of contractions. For instance, “isn’t" became “is not."
@@
ML Models Accuracy of Lexicon based methods
@@
To rigorously evaluate tweet sentiment classification, we employed a diverse set of DNN models, each with distinct architectural characteristics. Our initial model, the ‘Single-Dense Layered Neural Network,’ started with a transformer for feature extraction. A single dense layer with 512 units and ReLU activation captured high-level representations. The simplicity of this architecture allowed us to establish a baseline for performance comparison.
@@
We introduced the ‘BiLSTM + 3 Hidden Dense Layers’ model to explore the nuances of DL further. This architecture incorporated a bidirectional Long Short-Term Memory (BiLSTM) layer, enabling the network to capture sequential dependencies in the input data. Following the BiLSTM layer, three additional dense layers (512, 256, and 128 units) were used to distill the features further. Dropout was applied to enhance model generalization.
@@
Our systematic exploration of these DNN architectures, coupled with a comprehensive array of text representations, provided valuable insights into their performance and allowed us to tailor our approach to the intricacies of tweet sentiment analysis. This diversity of models served as a critical asset in our research, enabling us to select the most effective strategies for this challenging task.
@@
The proposed TRABSA model presents a systematic and effective architecture for sentiment analysis, as shown in Figure 5. The model utilizes the ‘cardiffnlp/twitter-roberta-base-sentiment-latest’ pre-trained transformer, capitalizing on its contextual understanding of the text. The architecture begins with input layers, including ‘input_ids’ and ‘attention_mask,’ where a maximum sequence length of 256 tokens is utilized. RoBERTa is used for its excellent performance in tasks involving natural language understanding. It encodes the input tweet text and generates contextual embeddings.
@@
The input data is then organized into a TensorFlow dataset, shuffled, and batched for training. The batch size is set to 16 to facilitate efficient training. The labels are one-hot encoded to prepare them for multiclass classification. The dataset is split into training and validation sets for rigorous evaluation. The core architecture of the TRABSA model is based on the Twitter RoBERTa architecture with specific enhancements:
@@
  • BiLSTM: A BiLSTM layer is introduced to capture contextual information from the text. The bidirectional nature of this layer allows the model to consider both past and future tokens in the sequence.
@@
The model is compiled using the Adam optimizer with a learning rate of $4\times 10^{-5}$ and categorical cross-entropy loss. Categorical accuracy is used as the evaluation metric. We included model checkpointing, early stopping, and callbacks to optimize model training. The early stopping mechanism monitors validation loss and restores the best weights to prevent overfitting. The training process involves fitting the model on the training dataset and validating it on the validation dataset for 50 epochs.
@@
Figure 7 comprehensively evaluates the TRABSA model’s performance in tweet sentiment analysis. The top left corner showcases the model’s classification metrics plot, illustrating precision, recall, and F1-score metrics for each sentiment class. Moving clockwise, the confusion matrix provides a detailed analysis of the model’s classification performance by comparing predicted sentiment labels with actual labels. The loss vs. epoch curve illustrates its training and validation loss over successive epochs, while the training and validation accuracy vs. epoch curve depicts the model’s learning progress and convergence. Together, these visualizations offer insights into the TRABSA model’s classification accuracy, convergence, and optimization process, aiding researchers in assessing its performance and identifying areas for improvement.
The TRABSA model’s architecture combines the strengths of RoBERTa’s contextual embeddings, BiLSTM’s sequence modeling, attention mechanisms for capturing interdependencies, and a well-designed set of dense layers to improve sentiment analysis accuracy for tweet data. This architecture demonstrated superior performance compared to other models, making it a noteworthy addition to the field of sentiment analysis.
@@ Results
Analyzing the performance of various models across different word embedding techniques reveals significant variations in their effectiveness for sentiment analysis tasks. Among the traditional models, GBM and LightGBM consistently demonstrate strong performance across all word embedding techniques, achieving macro average F1-scores ranging from 79% to 83%. Pre-trained transformer models like BERT and RoBERTa exhibit competitive performance, with the RoBERTa model showcasing high accuracy across various configurations, especially when combined with advanced neural network architectures. Notably, the TRABSA model outperforms all other models across all evaluation metrics, demonstrating exceptional macro average precision (94%), macro average recall (93%), macro average F1-score (94%), and accuracy of 94% (see Table LABEL:tab3). This significant improvement underscores the efficacy of the TRABSA framework in sentiment analysis tasks, surpassing even state-of-the-art transformer models like RoBERTa.
@@
We used BoW, TFIDF, word2vec, BERT, SBERT, and RoBERTa as word embeddings with all the ML and two DL models. When we noticed a significant improvement in the accuracy (above 80%) while using RoBERTa word embeddings with simple DL architectures, we decided to implement more complex hybrid DL models: BiLSTM+3 Hidden Layers NN, BiLSTM+CNN, and TRABSA.
Table 3: Performance evaluation of various models and embeddings

### v3

# RoBERTa and Attention-based BiLSTM for Interpretable Sentiment Analysis of Tweets

@@ Introduction
This study addresses the need for robust and generalizable SA models by proposing the "Transformer-based Attention Bidirectional LSTM for Sentiment Analysis (TRABSA)" model. The TRABSA model integrates the strengths of transformer-based architectures with attention mechanisms and bidirectional long short-term memory (BiLSTM) networks to enhance the performance and adaptability of SA tasks. Our method seeks to capture the broad diversity of English language usage and offer a more thorough knowledge of sentiment expression in various linguistic situations by combining data from several locations into a single dataset. By using this method, TRABSA can more effectively adjust to the subtle differences in English language usage across various groups, which improves the precision and significance of SA findings. We test the performance of the TRABSA model on a range of DL architectures and datasets, including extensive Twitter and external social media datasets, with a particular emphasis on scalability, accuracy, and consistency. Additionally, we do interpretability assessments utilizing the SHAP and LIME approaches to understand the model’s decision-making mechanism. We show the TRABSA model’s generalizability and robustness through our thorough examination, providing a viable method for SA in various real-world situations.
@@
    We conduct extensive experiments to assess the TRABSA model’s performance on the UK COVID-19 Twitter Dataset, benchmarking against 7 traditional machine learning (ML) models, 4 stacking models, and 4 DL models, demonstrating its superiority and versatility.
@@
The rest of this article is structured as follows: Section "Related Works" reviews state-of-the-art literature in SA using ML-DL and interpretability techniques. In the section "Methodology," the data collection and preprocessing techniques, unsupervised text labeling, implemented ML and DL models for benchmarking, architecture, and methodology of the proposed TRABSA model are described. The section "Results" presents the experimental setup, evaluation metrics, results, and robustness analysis of the TRABSA model. The SHAP and LIME analysis conducted on the TRABSA model are covered in section "Interpretability Analysis". Section "Disucssions" addresses the findings and implications of our investigation. In conclusion, the article is summarized, and future research directions are outlined in section "Conclusions and Future Directions".
@@
Recent advancements in SA and event detection have introduced several innovative models. DocTopic2Vec, proposed by Truică et al.[24], enhances document-level SA by combining local and global contexts through document and topic embeddings, outperforming traditional methods. EDSA-Ensemble[25] improves sentiment classification on social media by integrating event detection with SA using an ensemble approach. Petrescu et al.[26] bridges network and content analysis by combining event detection with SA, achieving high accuracy in sentiment determination. For imbalanced datasets, Truică and Leordeanu[27] compare machine learning algorithms, emphasizing the impact of dataset characteristics on classification performance. Lastly, ATESA-BÆRT by Apostol et al.[28] addresses aspect-based SA using a transformers-based ensemble, outperforming existing models in handling reviews with multiple aspects. Additionally, Mitroi et al.[29] introduces TOPICDOC2VEC, a new topic-document embedding that combines DOC2VEC and TOPIC2VEC, showing superior performance in polarity detection using game reviews.
@@
Our proposed methodological framework outlines a structured approach to SA, encompassing several key stages to ensure robustness and effectiveness in model development and evaluation, as shown in Figure 1. The first stage involves gathering relevant data for SA. We extend existing datasets by collecting additional tweets from diverse sources to enhance the dataset’s representativeness and coverage. Following data collection, we perform cleaning and preprocessing tasks to ensure the quality and consistency of the data. This includes removing noise, expanding contractions, handling duplicates, emojis, and missing tweets, and standardizing text formats. The next step involves labeling the data into positive, neutral, and negative sentiments. We leverage the latest updated RoBERTa-based pre-trained transformer model for tokenization and sentiment labeling, enabling accurate and efficient text data processing. The dataset is divided into training, validation, and test sets after it has been labeled. This enables us to use the test set to assess the model’s performance on untested data, refine hyperparameters using the validation set, and train the model on a portion of the data. In this step, we build the SA model based on our suggested hybrid DL architecture. We used Keras Tuner for hyperparameter optimization within the specified search space, focusing on minimizing the validation loss as our objective to achieve the best performance. Following model development, we compare the trained model’s performance against baseline models or current state-of-the-art methods. We examine the model’s ability to correctly classify sentiments into positive, neutral, and negative categories using a variety of metrics, including accuracy, precision, recall, and F1-score. Finally, we employ XAI techniques to interpret the model’s predictions and gain insights into its decision-making process. This involves analyzing the model’s internal mechanisms, such as attention weights or feature importance, to understand the factors influencing its predictions and enhance model interpretability.
@@
In this section, we provide a comprehensive overview of the data collection and preprocessing procedures undertaken in our study, which laid the foundation for robust SA of tweets.
@@
Extended Datasets: To augment our research’s cross-cultural dimension and overcome the geographic limitations of the benchmark dataset proposed by Qi and Shabrina, we extended the existing UK COVID-19 Twitter dataset [16]. The tweets were sourced through a combination of data extraction tools, specifically Twint and the Twitter Academic API. These tools were chosen because of their ability to acquire tweets with geolocation information, which is crucial for conducting geographical analyses. However, it should be noted that only a few 1% of Twitter users actively opt to share their geographic location when composing tweets, and this feature is not enabled by default [46]. The extended datasets, comprising the Global Twitter COVID-19 Dataset and the USA Twitter COVID-19 Dataset, are publicly available in the Extended Covid Twitter Datasets repository [47].
To ensure a comprehensive dataset, we merged the data collected by Twint and the Twitter Academic API. This amalgamation allowed us to access a larger volume of tweets. In identifying tweets related to the COVID-19 pandemic, we employed specific keywords such as "corona" or "covid" in the Twint search configurations and the query field of the Twitter Academic API. This search strategy enabled us to extract tweets and associated hashtags containing these pertinent terms.
@@
    Extended Global COVID-19 Dataset: This extension involved the comprehensive scraping of 411,885 tweets from 32 English-speaking countries. This dataset expansion allowed us to capture sentiment variations across diverse English-speaking regions, as illustrated in Figure 3. In particular, cities such as the "United States," the "United Kingdom," "Australia," and "New Zealand" exhibit high tweet volumes, while several other cities have comparatively lower tweet counts. Figure 4 illustrates word clouds and word frequencies within tweets of the extended datasets. Figure 4 (left) represents a visual summary of the most frequently occurring words in a vast dataset related to the COVID-19 pandemic. At the center of this cloud is the word "covid," which dominates with a staggering 226,463 mentions. Other significant terms around it, such as "vaccine," "case," "test," and "people," indicate the key topics and concerns worldwide during the pandemic. Words like "death," "pandemic," and "health" also hold prominence, highlighting the gravity of public health issues. Additionally, terms like "Trump" and "government" suggest the political dimensions entwined with the pandemic discourse.
@@
    Extended USA COVID-19 Dataset: In addition to the international extension, we further enriched our data by creating an extended dataset focusing exclusively on the United States. This dataset comprised 7,500 tweets meticulously scraped from U.S.-based sources. Including this dataset allows for a closer examination of sentiment dynamics within a specific geographical context. Figure 4 (right) specific to the United States reveals notable trends and sentiments within the country during the COVID-19 pandemic. In this dataset, words like "corona," "coronavirus," and "covid" are prominent, underlining the ubiquitous presence of these terms in American discussions. Interestingly, words like "fool" and "joke" appear, possibly reflecting a spectrum of attitudes towards the pandemic response. Negative expressions like "shit," "fuck," and "die" also emerge, suggesting the emotional intensity and frustration associated with the situation. Terms like "test" and "case" point towards testing and infection rates concerns, while "april" indicates a temporal reference.
Figure 4: This word cloud visualizes the most frequently occurring words in the "Global" (left) and "Only USA" (right) datasets of COVID-19-related tweets. The size of each word corresponds to its frequency in the dataset.
@@
    Removal of Irrelevant Elements: We methodically eliminated any superfluous content that has no bearing on SA, including hashtags (#subject), stated usernames (@username), and any hyperlinks beginning with "www," "http," or "https." We also removed terms that were less than two characters and stop words. Stop words, though common in text, often lack significant sentiment polarity. Despite being often used in texts, stop words frequently lack strong emotive polarity. It’s important to note that negations like "not" and "no" were kept in because removing them may change the sense of whole sentences.
@@
    Handling Repeated Characters: Some users utilize repeating characters in their tweets to highlight intense feelings. Words not present in standard lexicons were transformed into their correct forms to standardize such expressions. For instance, "sooooo goooood" was normalized to "so good."
@@
    Extending Contractions: Removing punctuation after a contraction, such "isn’t" or "don’t," presented difficulties. They were expanded into their full forms to maintain the meaningfulness of contractions. For instance, "isn’t" became "is not."
@@
ML Model Accuracy of Lexicon based methods
@@
To rigorously evaluate tweet sentiment classification, we employed a diverse set of DNN models, each with distinct architectural characteristics. We utilized the Keras Tuner for hyperparameter tuning across these models to ensure optimal performance. The search space included LSTM layer units ranging from 128 to 768, dense layer units from 64 to 512, dropout rates from 0.1 to 0.5, and learning rates from $1\times 10^{-5}$ to $1\times 10^{-2}$. Using Keras Tuner’s Random Search method, we identified the best parameters for each model, significantly enhancing their performance. This systematic exploration, coupled with comprehensive text representations and optimized hyperparameters, provided valuable insights into the performance of various DNN architectures.
Our initial model, the ‘Single-Dense Layered Neural Network,’ started with a transformer for feature extraction. A single dense layer with 512 units and ReLU activation captured high-level representations. The simplicity of this architecture allowed us to establish a baseline for performance comparison.
@@
We introduced the ‘BiLSTM + 3 Hidden Dense Layers’ model to explore the nuances of DL further. This architecture incorporated a BiLSTM layer, enabling the network to capture sequential dependencies in the input data. Following the BiLSTM layer, three additional dense layers (512, 256, and 128 units) were used to distill the features further. Dropout was applied to enhance model generalization.
@@
The proposed TRABSA model presents a systematic and effective architecture for SA, as shown in Figure 5. The model utilizes the ‘cardiffnlp/twitter-roberta-base-sentiment-latest’ pre-trained transformer, capitalizing on its contextual understanding of the text. The architecture begins with input layers, including ‘input_ids’ and ‘attention_mask,’ where a maximum sequence length of 256 tokens is utilized. RoBERTa is used for its excellent performance in tasks involving natural language understanding. It encodes the input tweet text and generates contextual embeddings.
@@
The input data is then organized into a TensorFlow dataset, shuffled, and batched for training. The batch size is set to 16 to facilitate efficient training. The labels are one-hot encoded to prepare them for multiclass classification. The dataset is split into training and validation sets for rigorous evaluation. In optimizing the TRABSA model, Keras Tuner plays a crucial role by systematically exploring a well-defined search space to fine-tune various hyperparameters. The search space includes the number of units in the BiLSTM layer, ranging from 128 to 512, which balances model complexity and efficiency. Additionally, three Dense layers are tuned with units ranging from 128 to 768, 64 to 512, and 32 to 256, respectively, affecting the model’s capacity and computational demands. The dropout rate varies from 0.1 to 0.5 to prevent overfitting, while the learning rate is explored logarithmically between 1e-5 and 1e-3 to optimize convergence speed. The tuning process employs Random Search to sample various hyperparameter combinations, providing an efficient way to explore the space without exhaustive search. After running multiple trials, the tuner identifies the best hyperparameter settings (see Figure 5) based on validation loss, ensuring an optimized balance between performance and computational efficiency.
The core architecture of the TRABSA model is based on the pre-trained RoBERTa-base model architecture with specific enhancements:
@@
  • BiLSTM: A BiLSTM layer with $512\times 2$ units is introduced to capture contextual information from the text. The bidirectional nature of this layer allows the model to consider both past and future tokens in the sequence.
@@
The model is compiled using the Adam optimizer with a learning rate of $4\times 10^{-5}$ and categorical cross-entropy loss. Categorical accuracy is used as the evaluation metric. We included model checkpointing, early stopping, and callbacks to optimize model training. The early stopping mechanism monitors validation loss and restores the best weights to prevent overfitting. The training process involves fitting the model on the training dataset and validating it on the validation dataset for 50 epochs; however, due to the early stopping mechanism, the iterations stop after 23 epochs.
@@
Figure 7 comprehensively evaluates the TRABSA model’s performance in tweet SA. The top left corner showcases the model’s classification metrics plot, illustrating precision, recall, and F1-score metrics for each sentiment class. Moving clockwise, the confusion matrix provides a detailed analysis of the model’s classification performance by comparing predicted sentiment labels with actual labels. The loss vs. epoch curve illustrates its training and validation loss over successive epochs, while the training and validation accuracy vs. epoch curve depicts the model’s learning progress and convergence. Together, these visualizations offer insights into the TRABSA model’s classification accuracy, convergence, and optimization process, aiding researchers in assessing its performance and identifying areas for improvement.
The TRABSA model’s architecture combines the strengths of RoBERTa’s contextual embeddings, BiLSTM’s sequence modeling, attention mechanisms for capturing interdependencies, and a well-designed set of dense layers to improve SA accuracy for tweet data. This architecture demonstrated superior performance compared to other models, making it a noteworthy addition to the field of SA.
@@ Results
Analyzing the performance of various models across different word embedding techniques reveals significant variations in their effectiveness for SA tasks. Among the traditional models, GBM and LightGBM consistently demonstrate strong performance across all word embedding techniques, achieving macro average F1-scores ranging from 79% to 83%. Pre-trained transformer models like BERT and RoBERTa exhibit competitive performance, with the RoBERTa model showcasing high accuracy across various configurations, especially when combined with advanced neural network architectures. Notably, the TRABSA model outperforms all other models across all evaluation metrics, demonstrating exceptional macro average precision (94%), macro average recall (93%), macro average F1-score (94%), and accuracy of 94% (see Table LABEL:tab3). This significant improvement underscores the efficacy of the TRABSA framework in SA tasks, surpassing even state-of-the-art transformer models like RoBERTa.
@@
We used BoW, TFIDF, word2vec, BERT, SBERT, and RoBERTa as word embeddings with several ML and two DL models. When we noticed a significant improvement in the accuracy (above 80%) while using RoBERTa word embeddings with simple DL architectures, we decided to implement more complex hybrid DL models: BiLSTM+3 Hidden Layers NN, BiLSTM+CNN, and TRABSA.
Table 3: Comprehensive 10-fold cross-validated mean performance evaluation metrics with standard deviations for various models and embeddings, including time performance (training and inference times)